solve for x
8(x-3)-5(x-2)= 6
8(2x-3)-5(3x-4)=10

Answer by nahataba(3)

1. first you would distribute.
2. 8(x-3)-5(x-2)=6 so you would do 8 times x and 8 times -3 which is 8x-24.
3. then you do -5 times x and -5 times -3 which is -5x+10 because two negatives equal a positive.
4. so you have 8x-24-5x+10=6
5. you then combine like terms. 8x+-5x=3x and -24+10=-14 so you get 3x-14=6.
6. you then you add positive 14 to both sides (you switch the signs) and get 3x=20.
7. you then divide both sides by 3 and get x=6 2/3 (six and two-thirds)

for the next one you follow the same steps.
1. 8 times 2x and 8 times -3 which is 16x-24.
2. -5 times 3x and -5 times -4 which is -15x+20.
3. 16x-24-15x+20=10
4. combine 16x and -15x which is 1x or just x and combine -24 and 20 which is 4.
5. x-4=10
6. add 4 to both sides.
7. x=14
you can check your answers by plugging the number in for x.